# Paseo de la Alameda 7, 46010, Valencia, Spain

### v1

@@ Abstract
The Heisenberg uncertainty principle is one of the fundamental pillars of quantum mechanics and quantum field theory. It is normally introduced by postulating the commutation relations $[\hat{x}^{i},\hat{p}^{j}]=i\hbar\delta^{ij}$. However, as suggested by some quantum gravity models and string theory, this basic principle no longer holds true in the presence of a minimal length, possible the Plank length, and modifications of the commutation have been proposed i.e., of the form $[\hat{x}^{\mu},\hat{p}^{\nu}]=-i\hbar(1+\beta_{0}\,\hat{p}^{2}/\Lambda^{2})\eta^{\mu\nu}$(plus possible additional terms). In this work we will consider the previous modified uncertainty principle in terms of an effective field theory, comment upon some theoretical subtleties that are often overlooked in the literature, and constrain , for the first time, the $\Lambda$ scale with the Compton high-energy experimental data. Our findings suggest that high-energy experiments are potentially sensitive to these corrections and could serve as an effective framework for probing possible violations of the Heisenberg uncertainty principle.
@@ 1 Introduction
The Generalized Uncertainty Principle (GUP) was first introduced by Maggiore:1993kv ; Kempf_1995 ; Scardigli_1999 and extensively analyzed in later studies Bang:2006va ; Bosso:2023aht ; Pedram:2011aa ; Galan:2007ns ; Das:2012rv ; Pedram:2011gw ; Mignemi:2011wh ; Nozari:2012gd ; Tedesco:2011iv ; Balasubramanian:2014pba ; Tkachuk:2013lta ; Bosso:2022vlz ; Pramanik:2013zy ; Bosso:2018uus ; Chung:2018btu ; Bosso:2020fos ; Bosso:2020jay ; Scardigli_1999 ; Hossenfelder_2006 ; Das_2016 ; Maccone_2014 ; Tawfik_2014 ; Hossenfelder_2013 ; Nenmeli_2021 ; Chen_2013 ; Basilakos_2010 ; Gecim_2017 ; Gecim_2018 ; KONISHI1990276 ; Chang_2011 ; Gomes_2022 ; Das_2008 ; ali ; Marin . It normally appears in quantum gravity models Scardigli_1999 ; Hossenfelder_2013 ; Basilakos_2010 ; Gecim_2018 ; Das_2008 ; ali ; Marin and string theory KONISHI1990276 ; Chang_2011 ; Hossenfelder:2012jw and is a consequence of the presence of a minimal length, which is usually associated with the Planck length. The GUP is mathematically introduced by modifying the standard commutation relations with additional terms, i.e.,
@@
with dimensionless constants $\beta_{0}$ and $\gamma_{0}$, and $\Lambda$, the energy scale where the GUP becomes relevant. As mentioned previously, $\Lambda$ is normally considered on the Planck scale, and therefore all current high-energy and precision experiments are, at least in principle, insensitive to the presence of these new terms. However, it is legitimate to consider extensions of the Heisenberg principle regardless of its original motivation from quantum gravity or string theory. Even if it constitutes one of the building blocks of modern physics, it has not yet been probed with high-energy experiments within an effective field theory framework, which is the main goal of our paper.
@@ 2.1 Minimal GUP
The minimally modified commutation relations from the expression (4), that respect (5), meaning that they do not give rise to non-commutative geometries that suffer from further complications Quesne:2006is ; Todorinov:2018arx , are given by
$[\hat{x}^{\mu},\hat{p}^{\nu}]=-i\eta^{\mu\nu}(1+\beta\hat{p}^{2})\,,$ (8)
@@ 2.2 Subtleties in building a modified effective Lagrangian and observables
$\qquad+\lambda\phi^{\dagger}\phi\Big{[}\partial^{\mu}(1-\beta\,\partial^{\nu}\partial_{\nu})\phi^{\dagger}\Big{]}\Big{[}\partial_{\mu}(1-\beta\,\partial^{\alpha}\partial_{\alpha})\phi\Big{]}\,.$ (15)
@@
$\big{<}\vec{p}_{3},\vec{p}_{4}\big{|}\mathcal{L}_{I}(x)\big{|}\vec{p}_{1},\vec{p}_{2}\big{>}=\big{<}0\big{|}a(\vec{p}_{3})\,b(\vec{p}_{4})\,\mathcal{L}_{I}(x)\,a^{\dagger}(\vec{p}_{1})\,b^{\dagger}(\vec{p}_{2})\big{|}0\big{>}\,.$ (17)
The subtlety, that is not often mentioned in the literature, consists in noticing that the ladder operators were quantized in terms of momenta $k^{\mu}$, that do dot obey momentum conservation. However, as these operators end up depending on physical momenta, four-momentum properties and conservation will hold as usual. In consequence, the terms in the transition matrix (17) will also depend on the physical momenta, an in general, all observables are defined in a physically meaningful way. The same conclusions can be reached using the path-integral approach, and even more clearly, as it eliminates the need to rely on the commutation relations.
@@ 3 Modified QED Lagrangian
$\mathcal{L}^{\text{GUP}}_{0}=\bar{\psi}\Big{(}i\gamma^{\mu}\partial_{\mu}(1-\beta\square)-m\Big{)}\psi-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}+\frac{\beta}{2}F^{\mu\nu}\Big{(}\square F_{\mu\nu}\Big{)}+\mathcal{O}(\beta^{2})\,,$ (19)
@@
∂̸$\not{\partial}\equiv\gamma^{\mu}\partial_{\mu}\,,\qquad\qquad\square\equiv\partial^{\alpha}\partial_{\alpha}\,.∂$ (20)
@@
$\mathcal{L}^{\text{GUP}}_{\text{QED}}$ $=\bar{\psi}\Big{(}i\gamma^{\mu}\partial_{\mu}(1-\beta\square)-m\Big{)}\psi-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}+\frac{\beta}{2}F^{\mu\nu}\Big{(}\square F_{\mu\nu}\Big{)}$
$\qquad\qquad-eQA_{\mu}\bar{\psi}\gamma^{\mu}\psi+\tilde{\mathcal{L}}_{I}^{\beta}+\mathcal{O}(\beta^{2})\,.$ (22)
where ~$\tilde{\mathcal{L}}_{I}^{\beta}$ is the $\mathcal{O}(\beta)$ interaction Lagrangian containing the newly generated interaction terms. It is given by
~$\tilde{\mathcal{L}}_{I}^{\beta}$ $=-i\beta\,\bar{\psi}\gamma^{\mu}D_{\mu}D^{\rho}D_{\rho}\psi\,.$ (23)
@@
$\psi\;\to\;\psi^{\prime}=e^{-ieQ\chi(x)}\psi\,,\qquad\qquad A_{\mu}\;\to\;A^{\prime}_{\mu}=A_{\mu}+\partial_{\mu}\chi(x)$ (24)
@@
~$\tilde{\mathcal{L}}_{I}^{\beta}$ $=\beta\,eQ\,\bar{\psi}\gamma^{\mu}\Big{(}A_{\mu}\square+2(\partial_{\mu}A^{\rho})\partial_{\rho}+2A^{\rho}\partial_{\rho}\partial_{\mu}+(\partial_{\mu}\partial_{\rho}A^{\rho})+(\partial_{\rho}A^{\rho})\partial_{\mu}\Big{)}\psi$
$\qquad+i\beta\,(eQ)^{2}\,\bar{\psi}\gamma^{\mu}\Big{(}A^{\rho}A_{\rho}\partial_{\mu}+2A_{\mu}A^{\rho}\partial_{\rho}+2A_{\rho}(\partial_{\mu}A^{\rho})+A_{\mu}(\partial^{\rho}A_{\rho})\Big{)}\psi$ (25)
$\qquad\qquad-\beta\,(eQ)^{3}\,\bar{\psi}\gamma^{\mu}A_{\mu}A^{\rho}A_{\rho}\psi\,,$
similar to the results found in Bosso:2020fos . [^2] Again, we found a sign difference with respect to Bosso:2020fos for the $(eQ)^{2}$ term, and extra terms, proportional to $(\partial_{\rho}A^{\rho})$ that are absent in the previously mentioned analysis. It is, therefore, necessary to point out that this type of terms do not cancel at the operator level. They are zero for on-shell photons only. One can trivially notice that the second line of the previous expression is not hermitian. However, we can easily obtain the hermitian version of the ~$\tilde{\mathcal{L}}_{I}^{\beta}$ interaction Lagrangian by summing the hermitian conjugate of the complex part (and multiply the sum by the corresponding 1/2 factor).
Finally, the hermitic expression for the $\mathcal{O}(\beta)$ interaction Lagrangian, denoted by ${\mathcal{L}}_{I}^{\beta}$ is given by
${\mathcal{L}}_{I}^{\beta}$ $=\beta\,eQ\,\bar{\psi}\gamma^{\mu}\Big{(}A_{\mu}\square+2(\partial_{\mu}A^{\rho})\partial_{\rho}+2A^{\rho}\partial_{\rho}\partial_{\mu}+(\partial_{\mu}\partial_{\rho}A^{\rho})+(\partial_{\rho}A^{\rho})\partial_{\mu}\Big{)}\psi$
$\qquad+\frac{i}{2}\beta\,(eQ)^{2}\,\bar{\psi}\gamma^{\mu}\Big{(}A^{\rho}A_{\rho}\partial_{\mu}-\overset{\leftarrow}{\partial}_{\mu}A^{\rho}A_{\rho}+2A_{\mu}A^{\rho}\partial_{\rho}-2\overset{\leftarrow}{\partial}_{\rho}A_{\mu}A^{\rho}\Big{)}\psi$ (26)
$\qquad\qquad-\beta\,(eQ)^{3}\,\bar{\psi}\gamma^{\mu}A_{\mu}A^{\rho}A_{\rho}\psi\,,$
Alternatively, one can initially part with the hermitic version of the Dirac Lagrangian.
$\mathcal{L}_{D}=\frac{i}{2}\bar{\psi}\gamma^{\mu}\overset{\leftrightarrow}{\partial}_{\mu}\psi-m\bar{\psi}\psi=\frac{i}{2}\left(\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi-(\partial_{\mu}\bar{\psi})\gamma^{\mu}\psi\right)-m\bar{\psi}\psi\,.$ (27)
After performing the $\partial_{\mu}\to\partial_{\mu}(1-\beta\partial^{\rho}\partial_{\rho})$ shift, and switching the ordinary partial derivative with the covariant derivative, one obtains the hermitic $\mathcal{O}(\beta)$ interaction Lagrangian.
@@
and not by $\overset{\leftarrow}{\partial}_{\mu}+ieQA_{\mu}$ in order to maintain gauge invariance.333 As a cross check, if this sign difference with respect to $D_{\mu}$ is not introduced, one does not obtain the standard QED interaction term. One can alternatively interpret this sign difference the following way: $D_{\mu}$ couples to $\psi$ with electric charge $Q_{\psi}=Q$, and $\overset{\leftarrow}{D}_{\mu}$ couples to $\bar{\psi}$ with the opposite electric charge $Q_{\bar{\psi}}=-Q$. However, one can check that the expressions (26) and (28) are equivalent i.e.,
@@ 3.1 Reduced interaction Lagrangian and Feynman rules
${\mathcal{L}}_{I}^{\beta}$ $=\beta\,eQ\,\bar{\psi}\gamma^{\mu}\Big{(}A_{\mu}\square+2(\partial_{\mu}A^{\rho})\partial_{\rho}+2A^{\rho}\partial_{\rho}\partial_{\mu}+(\partial_{\mu}\partial_{\rho}A^{\rho})+(\partial_{\rho}A^{\rho})\partial_{\mu}\Big{)}\psi$
$\qquad+\frac{i}{2}\beta\,(eQ)^{2}\,\bar{\psi}\gamma^{\mu}\Big{(}A^{\rho}A_{\rho}D_{\mu}-\overset{\leftarrow}{D}_{\mu}A^{\rho}A_{\rho}+2A_{\mu}A^{\rho}\partial_{\rho}-2\overset{\leftarrow}{\partial}_{\rho}A_{\mu}A^{\rho}\Big{)}\psi\,,$ (32)
@@
${\mathcal{L}}_{I}^{\beta}$ $=\beta\,eQ\,\bar{\psi}\gamma^{\mu}\Big{(}A_{\mu}\square+2(\partial_{\mu}A^{\rho})\partial_{\rho}+2A^{\rho}\partial_{\rho}\partial_{\mu}+(\partial_{\mu}\partial_{\rho}A^{\rho})+(\partial_{\rho}A^{\rho})\partial_{\mu}\Big{)}\psi$
$\qquad+m\,\beta\,(eQ)^{2}\,A^{\rho}A_{\rho}\bar{\psi}\psi+\frac{i}{2}\beta\,(eQ)^{2}\,\bar{\psi}\gamma^{\mu}\Big{(}2A_{\mu}A^{\rho}\partial_{\rho}-2\overset{\leftarrow}{\partial}_{\rho}A_{\mu}A^{\rho}\Big{)}\psi\,,$ (33)
@@
The derivation of the Feynman propagators for the SM fields (scalar, fermionic and gauge fields) can be found in Appendix A. Ignoring the $i\epsilon$ term, the Feynman rules for the modified propagators are shown in Figure 1, where we have split the full propagators (labeled as GUP) into the ordinary Standard Model contributions and, the new $\mathcal{O}(\beta)$ contributions. The new vertices at $\mathcal{O}(\beta)$ are given in Figure 2.444For on-shell photons, the term proportional to $p_{3}^{\mu}$, from the $\Gamma^{\mu}$ vertex, vanishes due to the orthogonality condition $\epsilon^{r}_{\mu}(p_{3})\,p_{3}^{\mu}=0$, where $\epsilon_{\mu}^{3}(p_{3})$ are the photon polarization vectors. Similarly, the terms proportional to $p_{1}^{\mu}$ and to $p_{3}^{\nu}$ also vanish for the $\Gamma^{\mu\nu}$ vertex.

### v2

# The Generalized Uncertainty Principle. New Bounds and Trends

@@ Abstract
The Heisenberg uncertainty principle is one of the fundamental pillars of quantum mechanics and quantum field theory. It is normally introduced by postulating the commutation relations $\hat{x}^{i},\hat{p}^{j}]=i\hbar\delta^{ij}$. However, as suggested by some quantum gravity models and string theory, this basic principle no longer holds true in the presence of a minimal length, possible the Plank length, and modifications of the commutation have been proposed i.e., of the form $\hat{x}^{\mu},\hat{p}^{\nu}]=-i\hbar(1+\beta_{0}\,\hat{p}^{2}/\Lambda^{2})\eta^{\mu\nu}$(plus possible additional terms). In this work we will consider the previous modified uncertainty principle in terms of an effective field theory, comment upon some theoretical subtleties that are often overlooked in the literature, and constrain, for the first time, the $\Lambda$ scale with the Compton high-energy experimental data. Our findings suggest that high-energy experiments are potentially sensitive to these corrections and could serve as an effective framework for probing possible violations of the Heisenberg uncertainty principle.
@@ 1 Introduction
The Generalized Uncertainty Principle (GUP) was first introduced by Maggiore:1993kv ; Kempf_1995 ; Scardigli_1999 and extensively analyzed in later studies Scardigli_1999 ; PhysRevD.108.066008 ; Bang:2006va ; Bosso:2023aht ; Pedram:2011aa ; Galan:2007ns ; Das:2012rv ; Pedram:2011gw ; Mignemi:2011wh ; Nozari:2012gd ; Tedesco:2011iv ; Balasubramanian:2014pba ; Tkachuk:2013lta ; Bosso:2022vlz ; Pramanik:2013zy ; Bosso:2018uus ; Chung:2018btu ; Bosso:2020fos ; Bosso:2020jay ; Hossenfelder_2006 ; Das_2016 ; Maccone_2014 ; Tawfik_2014 ; Hossenfelder_2013 ; Nenmeli_2021 ; Chen_2013 ; Basilakos_2010 ; Gecim_2017 ; Gecim_2018 ; KONISHI1990276 ; Chang_2011 ; Gomes_2022 ; Das_2008 ; ali ; Marin . It normally appears in quantum gravity models Scardigli_1999 ; Hossenfelder_2013 ; Basilakos_2010 ; Gecim_2018 ; Das_2008 ; ali ; Marin and string theory KONISHI1990276 ; Chang_2011 ; Hossenfelder:2012jw and is a consequence of the presence of a minimal length, which is usually associated with the Planck length. The GUP is mathematically introduced by modifying the standard commutation relations with additional terms, i.e.,
@@
with dimensionless constants $\beta_{0}$ and $\gamma_{0}$, and $\Lambda$, the energy scale where the GUP becomes relevant, and where $\hat{p}^{2}$ is given by the scalar product
$\hat{p}^{2}\equiv\hat{p}^{\mu}\hat{p}_{\mu}\,.$ (3)
As mentioned previously, $\Lambda$ is normally considered on the Planck scale, and therefore all current high-energy and precision experiments are, at least in principle, insensitive to the presence of these new terms. However, it is legitimate to consider extensions of the Heisenberg principle regardless of its original motivation from quantum gravity or string theory. Even if it constitutes one of the building blocks of modern physics, it has not yet been probed with high-energy experiments within an effective field theory framework, which is the main goal of our paper.
@@ 1.2 Minimal GUP
The minimally modified commutation relations from the expression (5), that respect (6), meaning that they do not give rise to non-commutative geometries that suffer from further complications Quesne:2006is ; Todorinov:2018arx , are given by (see Appendix A for further details)
$[\hat{x}^{\mu},\hat{p}^{\nu}]=-i\eta^{\mu\nu}(1+\beta\hat{p}^{2})-\,2i\beta\hat{p}^{\mu}\hat{p}^{\nu},$ (9)
@@ 1.3 Subtleties in building a modified effective Lagrangian and observables
$+\lambda\phi^{\dagger}\phi\Big[\partial^{\mu}(1-\beta\,\partial^{\nu}\partial_{\nu})\phi^{\dagger}\Big]\Big[\partial_{\mu}(1-\beta\,\partial^{\alpha}\partial_{\alpha})\phi\Big]\,.$ (16)
@@
$\big<\vec{p}_{3},\vec{p}_{4}\big|$ $\mathcal{L}_{I}(x)\big|\vec{p}_{1},\vec{p}_{2}\big$>=
$\big<0\big|a(\vec{p}_{3})\,b(\vec{p}_{4})\,\mathcal{L}_{I}(x)\,a^{\dagger}(\vec{p}_{1})\,b^{\dagger}(\vec{p}_{2})\big|0\big$>\,. (18)
The subtlety, that is not often mentioned in the literature, consists in noticing that the ladder operators were quantized in terms of momenta $k^{\mu}$, that do not obey momentum conservation. However, as these operators end up depending on physical momenta, four-momentum properties and conservation will hold as usual. In consequence, the terms in the transition matrix (1.3) will also depend on the physical momenta, an in general, all observables are defined in a physically meaningful way. The same conclusions can be reached using the path-integral approach, and even more clearly, as it eliminates the need to rely on the commutation relations.
@@ 2 Modified QED Lagrangian
$\mathcal{L}^{\text{GUP}}_{0}$ $=\bar{\psi}\Big(i\gamma^{\mu}\partial_{\mu}(1-\beta\square)-m\Big)\psi$
$\quad-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}+\frac{\beta}{2}F^{\mu\nu}\Big(\square F_{\mu\nu}\Big)+\mathcal{O}(\beta^{2})\,$, (20)
@@
$\not{\partial}\equiv\gamma^{\mu}\partial_{\mu}\,,\qquad\qquad\square\equiv\partial^{\alpha}\partial_{\alpha}\,.$ (21)
@@
$\mathcal{L}^{\text{GUP}}_{\text{QED}}$ $=\bar{\psi}\Big(i\gamma^{\mu}\partial_{\mu}(1-\beta\Box)-m\Big)\psi-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}$
$\quad+\frac{\beta}{2}F^{\mu\nu}(\Box F_{\mu\nu})-eQA_{\mu}\bar{\psi}\gamma^{\mu}\psi$
$\quad\quad+\tilde{\mathcal{L}}_{I}^{\beta}+\mathcal{O}(\beta^{2})\,.$ (23)
where $\tilde{\mathcal{L}}_{I}^{\beta}$ is the $\mathcal{O}(\beta)$ interaction Lagrangian containing the newly generated interaction terms. It is given by
$\tilde{\mathcal{L}}_{I}^{\beta}$ $=-i\beta\,\bar{\psi}\gamma^{\mu}D_{\mu}D^{\rho}D_{\rho}\psi\,.$ (24)
@@
$\psi\;\to\;\psi^{\prime}=e^{-ieQ\chi(x)}\psi\,$,
$A_{\mu}\;\to\;A^{\prime}_{\mu}=A_{\mu}+\partial_{\mu}\chi(x)\,$, (25)
@@
$\tilde{\mathcal{L}}_{I}^{\beta}$ $=\beta\,eQ\,\bar{\psi}\gamma^{\mu}\Big[A_{\mu}\Box+2(\partial_{\mu}A^{\rho})\,\partial_{\rho}+2A^{\rho}\,\partial_{\rho}\partial_{\mu}$
$\quad+(\partial_{\mu}\partial_{\rho}A^{\rho})+(\partial_{\rho}A^{\rho})\,\partial_{\mu}\Big]\psi$
$\quad\quad+i\beta\,(eQ)^{2}\,\bar{\psi}\gamma^{\mu}\Big[A^{\rho}A_{\rho}\,\partial_{\mu}+2A_{\mu}A^{\rho}\,\partial_{\rho}$
$\quad\quad\quad+2A_{\rho}(\partial_{\mu}A^{\rho})+A_{\mu}(\partial^{\rho}A_{\rho})\Big]\psi$
$\quad\quad\quad\quad-\beta\,(eQ)^{3}\,\bar{\psi}\gamma^{\mu}A_{\mu}A^{\rho}A_{\rho}\,\psi\,$, (26)
similar to the results found in Bosso:2020fos . Again, we found a sign difference with respect to Bosso:2020fos for the $(eQ)^{2}$ term, and extra terms, proportional to $(\partial_{\rho}A^{\rho})$ that are absent in the previously mentioned analysis. It is, therefore, necessary to point out that this type of terms do not cancel at the operator level. They are zero for on-shell photons only.
One can trivially notice that the second line of the previous expression is not Hermitian. However, we can easily obtain the Hermitian version of the $\tilde{\mathcal{L}}_{I}^{\beta}$ interaction Lagrangian by summing the Hermitian conjugate of the complex part (and multiply the sum by the corresponding 1/2 factor).
Finally, the Hermitian expression for the $\mathcal{O}(\beta)$ interaction Lagrangian, denoted by ${\mathcal{L}}_{I}^{\beta}$ is given by
${\mathcal{L}}_{I}^{\beta}$ $=\beta\,eQ\,\bar{\psi}\gamma^{\mu}\Big[A_{\mu}\Box+2(\partial_{\mu}A^{\rho})\,\partial_{\rho}+2A^{\rho}\,\partial_{\rho}\partial_{\mu}$
$\quad+(\partial_{\mu}\partial_{\rho}A^{\rho})+(\partial_{\rho}A^{\rho})\,\partial_{\mu}\Big]\psi$
$\quad\quad+\frac{i}{2}\beta\,(eQ)^{2}\,\bar{\psi}\gamma^{\mu}\Big[A^{\rho}A_{\rho}\,\partial_{\mu}-\overset{\leftarrow}{\partial}_{\mu}A^{\rho}A_{\rho}$
$\quad\quad\quad+2A_{\mu}A^{\rho}\,\partial_{\rho}-2\overset{\leftarrow}{\partial}_{\rho}A_{\mu}A^{\rho}\Big]\psi$
$\quad\quad\quad\quad-\beta\,(eQ)^{3}\,\bar{\psi}\gamma^{\mu}A_{\mu}A^{\rho}A_{\rho}\,\psi\,.$ (27)
Alternatively, one can initially part with the Hermitian version of the Dirac Lagrangian.
$\mathcal{L}_{D}$ $=\frac{i}{2}\bar{\psi}\gamma^{\mu}\overset{\leftrightarrow}{\partial}_{\mu}\psi-m\bar{\psi}\psi$
$=\frac{i}{2}\left(\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi-(\partial_{\mu}\bar{\psi})\gamma^{\mu}\psi\right)-m\bar{\psi}\psi\,.$ (28)
After performing the $\partial_{\mu}\to\partial_{\mu}(1-\beta\partial^{\rho}\partial_{\rho})$ shift, and switching the ordinary partial derivative with the covariant derivative, one obtains the Hermitian $\mathcal{O}(\beta)$ interaction Lagrangian.
@@
and not by $\overset{\leftarrow}{\partial}_{\mu}+ieQA_{\mu}$ in order to maintain gauge invariance. [^2] As a cross check, if this sign difference with respect to $D_{\mu}$ is not introduced, one does not obtain the standard QED interaction term. One can alternatively interpret this sign difference the following way: $D_{\mu}$ couples to $\psi$ with electric charge $Q_{\psi}=Q$, and $\overset{\leftarrow}{D}_{\mu}$ couples to $\bar{\psi}$ with the opposite electric charge $Q_{\bar{\psi}}=-Q$. However, one can check that the expressions (26) and (29) are equivalent i.e.,
@@ 2.1 Reduced interaction Lagrangian and Feynman rules
${\mathcal{L}}_{I}^{\beta}$ $=\beta\,eQ\,\bar{\psi}\gamma^{\mu}\Big[A_{\mu}\Box+2(\partial_{\mu}A^{\rho})\,\partial_{\rho}+2A^{\rho}\,\partial_{\rho}\partial_{\mu}$
$\quad+(\partial_{\mu}\partial_{\rho}A^{\rho})+(\partial_{\rho}A^{\rho})\,\partial_{\mu}\Big]\psi$
$\quad\quad+\frac{i}{2}\beta\,(eQ)^{2}\,\bar{\psi}\gamma^{\mu}\Big[A^{\rho}A_{\rho}\,D_{\mu}-\overset{\leftarrow}{D}_{\mu}A^{\rho}A_{\rho}$
$\quad\quad\quad+2A_{\mu}A^{\rho}\,\partial_{\rho}-2\overset{\leftarrow}{\partial}_{\rho}A_{\mu}A^{\rho}\Big]\psi\,.$ (33)
@@
${\mathcal{L}}_{I}^{\beta}$ $=\beta\,eQ\,\bar{\psi}\gamma^{\mu}\Big[A_{\mu}\Box+2(\partial_{\mu}A^{\rho})\,\partial_{\rho}+2A^{\rho}\,\partial_{\rho}\partial_{\mu}$
$\quad+(\partial_{\mu}\partial_{\rho}A^{\rho})+(\partial_{\rho}A^{\rho})\,\partial_{\mu}\Big]\psi$
$\quad\quad+\frac{i}{2}\beta\,(eQ)^{2}\,\bar{\psi}\gamma^{\mu}\Big[2A_{\mu}A^{\rho}\,\partial_{\rho}-2\overset{\leftarrow}{\partial}_{\rho}A_{\mu}A^{\rho}\Big]\psi\,$
$\quad\quad\quad+m\,\beta\,(eQ)^{2}\,A^{\rho}A_{\rho}\,\bar{\psi}\psi.$ (34)
@@
The derivation of the Feynman propagators for the SM fields (scalar, fermionic and gauge fields) can be found in Appendix B. Ignoring the $i\epsilon$ term, the Feynman rules for the modified propagators are shown in Figure 1, where we have split the full propagators (labeled as GUP) into the ordinary Standard Model contributions and, the new $\mathcal{O}(\beta)$ contributions. The new vertices at $\mathcal{O}(\beta)$ are given in Figure 2.333For on-shell photons, the term proportional to $p_{3}^{\mu}$, from the $\Gamma^{\mu}$ vertex, vanishes due to the orthogonality condition $\epsilon^{r}_{\mu}(p_{3})\,p_{3}^{\mu}=0$, where $\epsilon_{\mu}^{3}(p_{3})$ are the photon polarization vectors. Similarly, the terms proportional to $p_{1}^{\mu}$ and to $p_{3}^{\nu}$ also vanish for the $\Gamma^{\mu\nu}$ vertex.